OpenStudy (anonymous):

lrene has a previous balance of $987 on a credit card with a 17.3% APR compounded monthly. If she made a payment of $50 this month, what is the new balance on her credit card? $1,157.75 $1,107.75 $937.00 $951.23

OpenStudy (tkhunny):

When did she make the payment? Beginning: (987 - 50) + (987 - 50)*(0.173/12) = 937 + 13.51 = 950.51 End: 987 + 987*(0.173/12) - 50 = 937 + 14.23 = 951.23 Some other time?

OpenStudy (anonymous):

thank you :)

OpenStudy (tkhunny):

Careful with these. Make your questioner specify when things happen. Don't let them make you guess. I may have time on an exam to work all the problem three times, but I'm guessing most humans, when first seeing the material, don't have that kind of time.

OpenStudy (anonymous):

All right, I'll make sure of that. I was confused on other questions like that, too. For example, one would ask something that had multiple answers (not a math question, but something pertaining to finances). Thank you!